OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists. 1/3 + 1/9 +1/27 + 1/38 ...

OpenStudy (anonymous):

s infry=a/1-r

OpenStudy (anonymous):

\[S \infty = a/1-r \]

OpenStudy (anonymous):

just find r

OpenStudy (anonymous):

yes now easy to find r

OpenStudy (anonymous):

is it different if r is greater than 1...

OpenStudy (anonymous):

if r>1 there is no infinite geometric series, it is only menat for r<1